JP launches a rocket straight upwards from a height of 20 feet above the ground with an initial upward velocity of 85 feet per second. The equation h(t) = -16tē + 85t + 20 models the height of the rocket off of the ground, t seconds after the rocket is launched.
a) Find the height of the rocket 3 seconds after it is launched.
b) At what time will the rocket reach the maximum height? What is the maximum height of the rocket? (4 pts)
c) When will the rocket reach a height of 100 feet?
d) At what time will the rocket hit the ground?
I know part a. is 131 feet. I need help with part b,c,and d. Please!

Answer by Alan3354(69443)

JP launches a rocket straight upwards from a height of 20 feet above the ground with an initial upward velocity of 85 feet per second. The equation h(t) = -16tē + 85t + 20 models the height of the rocket off of the ground, t seconds after the rocket is launched.
a) Find the height of the rocket 3 seconds after it is launched.
b) At what time will the rocket reach the maximum height?
Max height is the vertex of the parabola, at t = -b/2a
t = -85/-32 = 85/32 seconds.
---
What is the maximum height of the rocket? (4 pts)
Sub 85/32 in the equation.
----------------
c) When will the rocket reach a height of 100 feet?
h(t) = -16tē + 85t + 20 = 100
Solve for t
--------------------
d) At what time will the rocket hit the ground?
h(t) = -16tē + 85t + 20 = 0
Solve for t